Calculate the freezing point and the boiling point of a solution that contains 35 g of C12H22O11 in 370 g of water.


1
Expert's answer
2021-11-16T11:40:01-0500

Moles of "C_{12}H_{22}O_{11}=\frac{35}{342}=0.102moles"


Mass of water "= 370g = 0.37kg"


Molarity "= \frac{0.102}{0.37}=0.28M/kg"


Kf= 1.86°c


Depression = "0.28\u00d71.86=0.52\u00b0c"

Freezing point "=0-0.52=-0.52 \u00b0c"


Kb=0.52°c


Elevation = "0.28\u00d70.52=0.15"

Boiling point ="100+0.15=100.15\u00b0c"